the area of a rectangular field is 12,750 square yards. The width is 150 feet. What is the length of the field in feet.
Answer by bucky(2189)

You need to work this problem all in one set of units.
.
Since a square yard is 3 feet by 3 feet, a square yard contains 9 square feet.
.
So if you multiply the 12,750 square yards of the field by 9 square feet per square yard,
you find that the field contains 12,750 * 9 = 114,750 square feet.
.
But the area of the field is found by taking the product of its length (L) times its width (W)
to get:
.
L * W = Area
.
But you know that the area is 114,750 square feet and the width of the field is 150 feet.
Substitute these values into the equation and you get:
.
L * 150 = 114,750
.
Solve for L by dividing both sides by 150 to get:
.
L = 114,750/150 = 765 feet
.
So the length of the field is 765 feet.
.
You can check this by multiplying 765 feet by 150 feet then divide that product by 9 square feet
per square yard and see if the answer isn't the 12,750 square yards the problem said it should be.
If you do the math you will find out that it is 12,750 square yards, so the answer of 765 ft
for the length L is correct.
.
Hope this helps you to understand the problem.